OpenStudy (isabel☺):

a persons mass while dieting decreases by 5 %. exercise creates muscle and reduces fat, causing the person's density to increase by 2%. calculate the percent change in the persons volume.

OpenStudy (isabel☺):

@jim_thompson5910

OpenStudy (isabel☺):

@amistre64

OpenStudy (isabel☺):

@Callisto

OpenStudy (chmvijay):

density =mass /volume

OpenStudy (chmvijay):

0.4 is that right

OpenStudy (chmvijay):

sory iam wrong wait

OpenStudy (isabel☺):

how? what is mass and the volume

OpenStudy (chmvijay):

10 i think

OpenStudy (isabel☺):

can show me your solution? :)

OpenStudy (chmvijay):

sorry this is again wrong

OpenStudy (chmvijay):

density=mass/volume Mass decreases by 5% => new mass = (95/100)m Density increases by 2% => new density = (102/100)d So, new volume = new mass/ new density = (95/102)v Change in volume = v - (95/102)v = (7/102)v Percent change = (7/102) x 100 = 6.86 %

OpenStudy (chmvijay):

this is right LOL

OpenStudy (isabel☺):

lol i just asked the same question on yahoo and you answered the same... Thank you ☺☺☺

OpenStudy (chmvijay):

LOL:) what to do we have too search google is God now days haaa haaa haaaa any way it was nice problem i liked it :)

OpenStudy (isabel☺):

haha !! ur funny but srsly thanks! :)

OpenStudy (chmvijay):

haa haaa haaa ur welcome :)